Suppose you have a student loan of ​$50000 with an APR of 12​% for 40 years. a. What are your required monthly​ payments? The required monthly payment is ​$ ___ ​(Do not round until the final answer. Then round to the nearest cent as​ needed.)

If paid off in 20 years?

The percentage of the total paid that is interest is ?
Answer #1

for 40 years"

prinipal loan amount P = 50000
interest per period i= 0.01
number of period n= 480
Periodc monthly payment A = Pi/(1-(1+i)-n)= 504.25

(for 20 years, this would be : 550.54)

for 40 years percentage of the total paid that is interest is =(504.25*480-50000)*100/(504.25*480)= 79.34 %

( 20 years percentage of the total paid that is interest is =(550.54*240-50000)*100/(550.54*240)= 62.16 %)